if 4/a=a 2 /16 , then check whether a is rational or irrational number.

if 4/a=a2/16 , then check whether a is rational or irrational number.

Grade:9

1 Answers

Rituraj Tiwari
3 years ago
4/a = a^2/16

a^3 = 64

a= 4

then the no. is rational no.